A test on car braking reaction times for mean between 18 and 30 years old have produced a mean

and standard deviation of 0.610 sec. and 0.123 sec, respectively. When 40 male drivers of this age

group were randomly selected and tested for their breaking reaction times, a mean of 0.587

second came out. At the α = 0.10 level of significance, test the claim of the driving instructor that

his graduates had faster reaction times.


Solution:

I. Statement of the Hypothesis:

Ho :

H1 :

II. Statistical Test:


III. Level of Significance and Critical Value:


IV. Computed value of Z:


V. Conclusion:


1
Expert's answer
2022-01-10T05:16:52-0500

"\bar x" =0.610

"\sigma" =0.123

"\mu" =0.587

n=40

(a).the test hypothesis

Ho :"\mu" "\le" 0.587

Ha:"\mu" >0.587

(b).Statistical test

z-test

(c).Level of significance is 10% and the critical value is1.645


(d).computed value of z

"z=\frac{\bar x-\mu}{\frac{\sigma}{\sqrt n}}"


"z=\frac{0.610-0.587}{\frac{0.123}{\sqrt 40}}"


"z=1.18264"


(e).conclusion

since the computed z value,1.18264 is less than the critical value 1.645 ,we fail to reject the null hypothesis.